FINANCIAL LEVERAGE EFFECTS

Firms HL and LL are identical except for their financial leverage ratios and the interest rates they pay on debt. Each has $20 million in invested capital, has $4 million of EBIT, and is in the 40% federal-plus-state tax bracket. Firm HL, however, has a debt-to-capital ratio of 55% and pays 11% interest on its debt, whereas LL has a 25% debt-to-capital ratio and pays only 9% interest on its debt. Neither firm uses preferred stock in its capital structure.

  1. Calculate the return on invested capital (ROIC) for each firm. Round your answers to two decimal places.

    ROIC for firm LL is   %
    ROIC for firm HL is   %

  2. Calculate the rate of return on equity (ROE) for each firm. Round your answers to two decimal places.

    ROE for firm LL is    %
    ROE for firm HL is    %

  3. Observing that HL has a higher ROE, LL's treasurer is thinking of raising the debt-to-capital ratio from 25% to 60% even though that would increase LL's interest rate on all debt to 15%. Calculate the new ROE for LL. Round your answer to two decimal places.

    %
Answer #1

The Invested capital of both HL and LL is $20 million

Debt ratio of HL is 55% and Debt ratio of LL is 25% =  

Debt of HL= $11 million Debt of LL = $5 million

So, Equity of HL = Total capital - Debt (as there is no preferred stock)

= 20-11 = $ 9 million

So, Equity of LL = Total capital - Debt (as there is no preferred stock)

= 20-5 = $ 15 million

a) ROIC = Net operating profit after tax/ Invested Capital

= EBIT * (1- tax rate ) / Invested capital

As both firms have same EBIT, Tax rate and Invested capital  

ROIC (HL) = ROIC (LL) = 4 * (1-0.4)/ 20 = 0.12 = 12.00%

ROE is calculated as , ROE = Profit after tax / Equity

Profit after tax calculation is done as shown in table below

All figures in $ million
HL LL New LL
Invested Capital 20 20 20
- Debt 11 5 12
- Equity 9 15 8
EBIT 4 4 4
Interest 1.21 0.45 1.8
EBT 2.79 3.55 2.2
Tax 1.116 1.42 0.88
PAT 1.674 2.13 1.32

So, ROE (HL) =1.674/9 = 0.186 = 18.60%

ROE (LL) = 2.13/15 =0.142 = 14.20%

New ROE of LL (with changed capital structure) is also calculated in the table above

ROE (LL) = 1.32/8 = 0.1650 = 16.50%